OpenStudy (anonymous):

find the solution of 15y-y^2=0

zepdrix (zepdrix):

\[\Large\rm 15y-yy=0\]Take a y out of each term,\[\Large\rm y(\qquad ?\qquad)=0\]What are we left with in the brackets? :o Remember how to factor?

OpenStudy (ribhu):

take y common y(15-y) = 0 so now either y can be zero or 15-y should be zero. so y = 0, 15

OpenStudy (ribhu):

@mcnabkt your solution with explanation.

OpenStudy (anonymous):

\[15y-y^2=0 \\y(15-y)=0\\y=0 \vee y=15\]

OpenStudy (anonymous):

im lost....

OpenStudy (ribhu):

why?

OpenStudy (anonymous):

because math is like a puzzle to me

OpenStudy (anonymous):

helpppp

OpenStudy (kendricklamar2014):

Look at the attachment I put up there ^ @mcnabkt it shows the steps it takes

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

h

OpenStudy (anonymous):

e

OpenStudy (anonymous):

l

OpenStudy (anonymous):

p

OpenStudy (kendricklamar2014):

Look at this @mcnabkt:

TheSmartOne (thesmartone):

Hi there :)